OpenStudy (anonymous):

Solve for x: -3(x - 3) = 21

OpenStudy (anonymous):

-3x +9 = 21 -3x = 21 - 9 -3x = 12 x = -(12/3) x = -4

OpenStudy (jaredstone4):

If you factor the -3 into (x-3) you get -3x + 9 = 21. (It is a positive 9 because a negative times a negative equals a positive) -3x + 9 =21 Subtracting 9 from each side, you get: -3x = 12 Divide each side by -3 and you get: x = -4.